The cost of a TV set is Rs 18875 and the cost of 1 VCR is Rs 10550. Mr Anuj wants to buy 3 TV sets and 2 VCRs. He has Rs sixty thousand with him. How much more money does he need to buy them

Grade:12

1 Answers

SJ
3 years ago
Welcome to Askiitians
Cost of TV= Rs. 18875

Cost of VCR= Rs. 10550

3 TV sets= (18875x 3)

= 56625

2 VCRs cost= ( 10550x 2)

= 21100

Total cost= 56625+ 21100 = 77725

Step 2:

Mr. Anuj's total money= 60,000

Money needed more= 77,725- 60,000= 17725

He needs Rs. 17725 more.
Thanks